OpenStudy (anonymous):

Find the extrema for f(x)=0.5e^(-x^2)

OpenStudy (anonymous):

\[f(x)=0.5e ^{-x ^{2}}\]

OpenStudy (anonymous):

I got the derivative \[f'(x)=-e ^{-x ^{2}}\] But I never took college algebra and never learned about e. I don't know how to solve for x.

OpenStudy (jhannybean):

\[\large f(x) = 0.5e^{-x^2}\]\[\large f'(x)=(0.5)(-2x)e^{-x^2}\]\[\large f'(x) = -xe^{-x^2}\]

OpenStudy (anonymous):

Oh yeah. I missed an x typing.

OpenStudy (jhannybean):

Now you have to set \(f'(x) = -xe^{-x^2} =0\) and solve for x to get your critical points. Look up the extreme value theorem for further assistance. http://tutorial.math.lamar.edu/Classes/CalcI/AbsExtrema.aspx

OpenStudy (anonymous):

Yeah, I've read everything I can find on the extreme value theorem online and in my textbook. I understand it. I don't understand HOW to solve for x in that equation. I even know the answer. I just don't know how to do it. Thanks anyway.

OpenStudy (jhannybean):

Alright. So you've gotten your first derivative, and you've set it equal to 0, but now you see that theyre multiplied to eachother. Evaluate them separately. \[\large -x=0\]\[\large e^{-x^2} = 0\] solving for x = 0, divide both sides by -1, you'll find that x=0. Once you understand the domain of \(e^x\) you'll find that the function never passes through 0, it comes really close to it. (If you've got a graphing calculator you can graph the function \(e^{-x^2}\). So your only extrema in this case would be 0.

OpenStudy (anonymous):

Thank you!

OpenStudy (jhannybean):

I'd suggest reading up on your natural log functions to better understand the function e.

OpenStudy (anonymous):

I know. I've done that. My Calc teacher and my stats teacher have both tried to help me with that. I have a college algebra book sitting open next to me with that section open. It doesn't help. I'm old and it doesn't stick in my brain. :)

OpenStudy (jhannybean):

:) I had to practice a ton of problems with derivatives and integrals of e to see how the function worked.